# Controllable Traffic Simulation through LLM-Guided Hierarchical Chain-of-Thought Reasoning

### v1

@@ IV-B Guided generation
Chain-of-thought mechanism. Our goal is to convert user descriptions into cost functions using LLMs and guide the generation through the method outlined in III-C. Existing methods often struggle with processing and understanding complex descriptions. Our approach leverages a chain-of-thought (CoT) mechanism to systematically break down these events into hierarchical structures, and generate corresponding cost functions step-by-step, ensuring accurate interpretation and alignment of traffic participant trajectories.
@@
Figure 2: Utilizing a logic relation diagram and examining the hierarchical structure of traffic elements, we employ the chain-of-thought mechanism to thoroughly analyze and interpret these hierarchical behaviors for controllable generation.
@@
Figure 2 illustrates the detailed CoT process. First, recognize the events. For multiple events, break them down into several single events. Then, determine the number of agents involved. If multiple agents are involved, identify each agent’s behavior. Next, assess whether each behavior is map-related. If it is, design an algorithm to explore the road topology $E=\{e_{ij}\}$. Finally, generate cost functions. For map-related descriptions, Frenet-based cost functions are recommended. We will show that with this step-by-step thinking manner, the LLM can effectively understand and process a wider range of complex descriptions.
@@
Frenet-based cost functions. Inspired by the success of the Frenet frame in motion planning, we utilize it to construct cost functions for descriptions related to map information. The Frenet frame is shown on the left side of Figure 2. It describes the relative position of trajectories concerning the reference path by the longitudinal distance $s$ and the lateral offset $d$, which have clear geometric meanings, making them easy for LLMs to understand and utilize. LLMs will generate cost functions based on $s$ and $d$, enabling the control of both longitudinal and lateral directions. We will demonstrate that using these Frenet-based cost functions significantly enhances the LLMs’ understanding of spatial relationships within traffic scenarios, leading to a substantial improvement in their ability to generate accurate cost functions.
@@ V Experiments
Category Method ADE FDE JSD Collision Off Road Per Scene
(m) $\downarrow$ (m) $\downarrow$ (×$\times 10^{-2}$) $\downarrow$ (%) $\downarrow$ (%) $\downarrow$ Gen. Time (s) $\downarrow$
Auto-regressive Methods Actions-Only [36] $4.81$ $11.89$ $10.4$ $19.9$ $27.6$ $3.3$
Decision Transformer [37] $1.56$ $3.07$ $8.4$ $5.3$ $\mathbf{11.0}$ $20.7$
CtRL-Sim [9] $1.25$ $2.04$ $\underline{7.9}$ $5.3$ $\mathbf{11.0}$ $8.2$
Diffusion-based Methods CTG++ [15] $1.73$ $4.02$ $\mathbf{7.4}$ $5.9$ $15.0$ $44.0$
Ours (32 diffusion steps) $\mathbf{1.17}$ $\mathbf{1.93}$ $8.2$ $\underline{5.1}$ $\underline{12.8}$ 20.4
Ours (50 diffusion steps) $\underline{1.18}$ $\underline{1.97}$ $8.6$ $\mathbf{4.9}$ $13.2$ 29.1
TABLE I: Results of traffic simulation on WOMD. We bolded the optimal result and underlined the second best result.
@@ V-A Experiment Setup
Evaluation. We begin by conducting traffic simulation experiments, following the experimental settings of CtRL-Sim [9]. Using a 1.1-second history and road topology as input, we simulate the future trajectories over an 8-second horizon. We use the Final Displacement Error (FDE) and Average Displacement Error (ADE) to measure reconstruction error. We employ a distributional metric (JSD) to evaluate realism, defined as the mean of the Jensen-Shannon Distances computed on linear speed, angular speed, acceleration, and distance to the nearest vehicle between real and simulated scenes. Additionally, common-sense metrics like Collision rate and Offroad rate are used.
Given our flexible model design, which does not require encoding historical data, our model can handle input consisting solely of a single frame of agent state and road topology, which is consistent with the problem setting for scenario generation. Thus, we further perform experiments on traffic scenario generation to validate the model. In line with experimental settings in this area, we evaluate the model using scene collision rate (SCR) and the Maximum Mean Discrepancy (MMD) metrics for both distance to the nearest object features ($\mathrm{MMD_{o}}$) and distance to road edge features ($\mathrm{MMD_{r}}$).
Lastly, we conduct experiments on controllable scenario generation. We select a range of common or complex scenario descriptions, and evaluate whether the generated cost functions accurately reflect these descriptions. We also visualize the results to assess if the model correctly follows the guidance of the generated functions for each scenario.
Baselines. For traffic simulation, we compare our method against recent auto-regressive methods, including Actions-Only (a modified version of Trajeglish [36]), Decision Transformer [37], and CtRL-Sim [9]. We also include the popular diffusion-based method CTG++ [15]. For scenario generation, we compare our approach to TrafficGen [38] and LCTGen [39]. In the controllable generation experiments, we compare the expressiveness of our CoT mechanism combined with Frenet-frame-based cost functions against the CTG++ prompt method.
Implementation Details. We set the number of layers for all attention modules to 3. The number of agents is set to $N_{a}=24$, and the number of lane polylines to $N_{m}=200$. We use 32 sampling steps for diffusion sampling, with other hyperparameters following EDM [34]. For closed-loop simulation, we set the re-plan rate to 2 Hz. For controllable generation, we use GPT4 [40] for cost function generation.
(a) User query: vehicle 1 cuts in vehicle 2, while vehicle 3 turns right.
Figure 3: GPT solutions and visualizations for some cases.
V-B Traffic simulation results
The traffic simulation results are presented in Table I. In alignment with the experimental setup, we query the LLM with the instruction ”vehicle should reach $target\_point$ and avoid collisions,” and then the generated cost function is used to guide the model. Our model achieves the best reconstruction performance and ranks second in terms of common-sense metrics. However, the JSD metric is slightly inferior to the current state-of-the-art methods. A potential reason for this is that our method directly outputs future states, whereas other methods generate actions, which may lead to better dynamic feasibility. As for generation speed, our model demonstrates a significant advantage over diffusion-based CTG++ and is competitive to some auto-regressive methods.
Method Per Scene Generation Time (s)
w/o map reuse w map reuse
CTG++ 44.0 -
Ours (50 diffusion steps) 32.5 29.1
Ours (32 diffusion steps) 23.2 20.4
TABLE II: Comparison of genaration time with CTG++.
@@ V-B Traffic simulation results
TABLE III: Results of traffic scenario generation.
To further validate the efficiency of our design, we conduct an ablation experiment presented in Table II. In addition to our 32-diffusion-step version, we also evaluate our method using 50 diffusion steps to align with CTG++. By incorporating historical data through conditional sampling, our model eliminates the need for historical data encoders, resulting in a considerable speedup. Moreover, by leveraging relative coordinates to enable map reuse, the generation speed is further enhanced.
The results of traffic generation are shown in Table III. Our model outperforms others across all metrics, further confirming its ability to generate realistic traffic scenarios.
V-C Controllable simulation results
We test a range of descriptions including basic rules, complex interactions and abnormal behaviors as shown in Table IV. To assess the contribution of each component, we conduct an ablation study.
In terms of basic rules, all tested methods demonstrated effective performance by employing functions like Norm for target point and ReLU for speed limit. These results suggest that LLMs do not confine the generation of cost functions to recommended forms (e.g., projection or Frenet), but instead allow for diverse, adaptable cost function formulations.
When addressing interactions such as lane changes, all methods performed well. However, for more complex road topologies—such as navigating the rightmost lane and approaching the next intersection—CoT methods demonstrated superior performance due to their enhanced map exploration capabilities. In agent interactions like cut-ins, CoT effectively understands and processes them by recognizing that one vehicle continues straight while another performs a lane change into its path, while CTG++ struggles to understand them. Although CoT correctly understands these behaviors, projection-based methods still occasionally generate incorrect cost functions. In contrast, the Frenet-based approach consistently performs accurately, as its clear geometric interpretations further enhance the LLM’s understanding of the scenario.
Regarding abnormal behaviors, the CoT+Frenet approach demonstrates impressive performance. The LLM generates $sin(d)$ to simulate the vehicle swaying left and right within the lane, and controls $d$ to be positive, ensuring the vehicle stays on the left side of the road edge. The Frenet-based approach allows the LLM to express these abnormal behaviors with concise function forms, highlighting the expressiveness of the Frenet frame and its ability to enhance the LLM’s understanding of complex traffic scenarios. Additionally, this capability to generate abnormal behaviors is particularly valuable for simulating long-tail scenarios.
Category Description CTG++ CoT+ Projection CoT+ Frenet
Basic rules vehicle drives at a speed limit 10m/s
@@ V-C Controllable simulation results
Complex interactions vehicle makes a left-lane-change
@@
Abnormal behaviors vehicle drives in reverse
@@
TABLE IV: Results of controllable simulation.
@@
We further combine some descriptions then query the LLM, and visualize the generation results, as shown in Fig 3. The LLM effectively decomposed the complex descriptions and generated the correct forms of cost functions. The visualizations further demonstrate that the model accurately follows the guidance of the generated functions.
@@ VI Conclusion
In this paper, we proposed a novel CoT-enhanced traffic simulation model. Our framework is constructed on a diffusion-based generative model, which we have optimized for enhanced computational efficiency and controllability. Our method is capable of taking users’ descriptions as input, and generate a wide range of scenarios, including complex interactions and long-tail events. We hope our method will offer significant contributions to the field.

### v2

@@ IV-B Guided generation
Our goal is to convert user descriptions into cost functions using LLMs and guide the generation through the method outlined in III-C. Existing approaches often face challenges in processing and comprehending complex descriptions. To address this issue, we propose an LLM-based framework comprising a high-level understanding module and a low-level refinement module.
@@
Figure 2: We systematically design the chain-of-thought mechanism to thoroughly analyze and interpret the hierarchical behaviors in traffic scenes for controllable generation.
@@
High-Level Understanding Module. The high-level module employs a systematically designed Chain-of-Thought (CoT) mechanism to reason about the interaction graph and decompose complex events into manageable components. Figure 2 illustrates the detailed CoT process. First, the module identifies the events. For multiple events, it breaks them down into individual single events. Next, it determines the number of agents involved. If multiple agents are present, it identifies each agent’s behavior. Subsequently, it evaluates whether each behavior is map-related. If so, it designs an algorithm to explore the road topology $E=\{e_{ij}\}$. Finally, it generates cost functions. For map-related descriptions, Frenet-based cost functions are recommended. We demonstrate that this structured reasoning approach enables the LLM to effectively understand and process a broader range of complex descriptions.
Low-Level Refinement Module. While the high-level module excels at generating accurate cost functions for complex descriptions, the resulting cost functions may occasionally be suboptimal or challenging to control through diffusion guidance. To address this, we introduce the low-level refinement module to further refine the cost functions. This module takes the low-level generated trajectory coordinates as input, enabling the LLM to directly analyze the generated trajectory data. Additionally, the corresponding cost function values from the trajectory optimization process are provided, offering further insights for the LLM’s analysis. Using these inputs, the LLM evaluates whether the trajectory aligns with the given description, whether the cost function accurately reflects the desired behavior, and whether each loss term has appropriate hyperparameters. For example, the LLM can directly assess the correctness of the behavior from the trajectory coordinates and identify potential issues from unusually high cost values, enabling precise error analysis and correction. Through this design, we enhance the framework’s ability to understand complex traffic scenarios and dynamically adapt to varying traffic contexts.
@@
Frenet-based cost functions. Inspired by the success of the Frenet frame in motion planning, we utilize it to construct cost functions for descriptions related to map information. The Frenet frame is shown on the left side of Figure 2. It describes the relative position of trajectories concerning the reference path by the longitudinal distance $s$ and the lateral offset $d$. These parameters have clear geometric interpretations, making them intuitive for LLMs to understand and utilize. Furthermore, unlike methods that rely on target points, which can overly constrain the feasible trajectory space, our Frenet-based cost functions provide more reasonable constraints by naturally decomposing trajectories into longitudinal and lateral components. LLMs will generate cost functions based on $s$ and $d$, enabling the control of both longitudinal and lateral directions. We will demonstrate that these Frenet-based cost functions significantly improve the LLMs’ understanding of spatial relationships in traffic scenarios, leading to more accurate cost function generation and higher guidance success rates.
Method ADE FDE JSD Collision Off Road
(m) $\downarrow$ (m) $\downarrow$ (×$\times 10^{-2}$) $\downarrow$ (%) $\downarrow$ (%) $\downarrow$
Actions-Only [38] 4.814.814.81 11.8911.8911.89 10.410.410.4 19.919.919.9 27.627.627.6
DT [39] 1.561.561.56 3.073.073.07 8.48.48.4 $\underline{5.3}$ $\mathbf{11.0}$
CtRL-Sim [9] $\underline{1.25}$ $\underline{2.04}$ $\underline{7.9}$ $\underline{5.3}$ $\mathbf{11.0}$
CTG++ [15] 1.731.731.73 4.024.024.02 $\mathbf{7.4}$ 5.95.95.9 15.015.015.0
Ours $\mathbf{1.17}$ $\mathbf{1.93}$ 8.28.28.2 $\mathbf{5.1}$ $\underline{12.8}$
TABLE I: Results of conditional simulation. We bolded the optimal result and underlined the second best result.
@@ V-A Experiment Setup
Evaluation. We begin by evaluating our model design through experiments on two tasks. The first task is conditional simulation, where we follow the experimental setup of CtRL-Sim [9]. In this setup, the model is provided with a 1.1-second history and an endpoint, and it simulates future trajectories over an 8-second horizon. We use the Final Displacement Error (FDE) and Average Displacement Error (ADE) to measure reconstruction error. We employ a distributional metric (JSD) to evaluate realism, defined as the mean of the Jensen-Shannon Distances computed on linear speed, angular speed, acceleration, and distance to the nearest vehicle between real and simulated scenes. Additionally, common-sense metrics like Collision rate and Offroad rate are used.
The second task is unconditional generation, where only the current states are provided. We evaluate the model using scene collision rate (SCR) and the Maximum Mean Discrepancy (MMD) metrics for both distance to the nearest object features ($\mathrm{MMD_{o}}$) and distance to road edge features ($\mathrm{MMD_{r}}$).
Finally, we conduct experiments on guided simulation. First, we select a range of scenario descriptions and assess whether the generated cost functions accurately reflect these descriptions. Next, we further choose four types of events, select 20 scenarios for each type, and evaluate the generation success rate. We visualize the results to determine whether the model correctly adheres to the guidance provided by the generated cost functions in each scenario.
@@
TABLE II: Results of unconditional scenario generation.
Baselines. For conditional simulation, we compare our method against recent auto-regressive methods, including Actions-Only (a modified version of Trajeglish [38]), Decision Transformer [39], and CtRL-Sim [9]. We also include the popular diffusion-based method CTG++ [15]. For unconditional generation, we compare our approach to TrafficGen [40] and LCTGen [41]. In the guided generation experiments, we compare our proposed LLM-based framework in comparison to CTG++ [15] and SceneDiffuser [16].
Implementation Details. We set the number of layers for all attention modules to 3. The number of agents is set to $N_{a}=24$, and the number of lane polylines to $N_{m}=200$. We use 32 sampling steps for diffusion sampling, with other hyperparameters following EDM [35]. For closed-loop simulation, we set the re-plan rate to 2 Hz. For guided simulation, we use GPT-4o [42] for cost function generation.
Category  Description CTG++
Scene
Diffuser
Ours
w/o understanding
Ours
zero-shot CoT
Ours
w/o Frenet
Ours
Basic
rules
vehicle drives at a speed limit 10m/s
@@
Behavior and
interactions
vehicle makes a left-lane-change
@@
Abnormal
behaviors
vehicle drives in reverse
@@
TABLE III: Results of controllable simulation.
V-B Traffic simulation results
The conditional simulation results are presented in Table I. Our model achieves the best reconstruction performance, as measured by FDE and ADE, and ranks second in terms of common-sense metrics such as Collision Rate and Offroad Rate. However, the Jensen-Shannon Distance (JSD) metric is slightly inferior to the current state-of-the-art methods. A potential reason for this is that our method directly outputs future states, whereas other methods generate actions, which may lead to better dynamic feasibility.
The results of unconditional generation, shown in Table II, further demonstrate the effectiveness of our approach. Our model outperforms all baselines across metrics such as Scene Collision Rate (SCR) and Maximum Mean Discrepancy (MMD), confirming its ability to generate realistic traffic scenarios. These two experiments show that our flexible model design can handle various tasks and achieve competitive performance.
V-C Guided simulation results
Evaluation on generated cost functions. Firstly, we evaluate our method using a diverse set of scenario descriptions, including basic rules, common behaviors and interactions, and abnormal behaviors, as detailed in Table III. We analyze the generated cost functions to verify their alignment with the input descriptions. To validate the contribution of each component in our method, we conduct an ablation study with the following variations: (1) our method without the high-level understanding module, and (2) our method without the Frenet coordinate system. Additionally, to further demonstrate the effectiveness of our design, we replace our CoT with zero-shot CoT [43], where the LLM is instructed to ”think step by step” and reason about the traffic event independently, and compare the results.
Scenario Type
Scene
Diffuser
Ours
w/o Refine
Ours
w/ Refine
Cut in 0.45 0.65 0.85
Out of road - 0.70 0.90
Yield 0.40 0.70 0.85
Rightmost - 0.75 0.95
TABLE IV: Results of guidance success rate.
In terms of basic rules, all tested methods demonstrated effective performance by employing functions like Norm for target point and ReLU for speed limit. These results suggest that LLMs do not confine the generation of cost functions to recommended forms (e.g., Frenet), but instead allow for diverse, adaptable cost function formulations.
When addressing behaviors such as lane changes, all methods performed well. However, for more complex road topologies such as navigating the rightmost lane, methods with our designed CoT demonstrated superior performance due to their enhanced map exploration capabilities. In agent interactions like cut-ins, our understanding module effectively understands and processes them by recognizing that one vehicle continues straight while another performs a lane change into its path, while CTG++ struggles to understand them. Although correctly understanding these behaviors, methods without Frenet coordinates cannot sufficiently understand the spatial relationships thus generating wrong functions. In contrast, the Frenet-based approach performs accurately, as its clear geometric interpretations further enhance the LLM’s understanding of the scenario.
Regarding abnormal behaviors, our approach demonstrates impressive performance. The LLM generates $sin(d)$ to simulate the vehicle swaying left and right within the lane, and controls $d$ to be positive, ensuring the vehicle stays on the left side of the road edge. The Frenet-based approach allows the LLM to express these abnormal behaviors with concise function forms, highlighting the expressiveness of the Frenet frame and its ability to enhance the LLM’s understanding of complex traffic scenarios.
Through this experiment, we demonstrate that our method can effectively handle a broader range of scene descriptions compared to current approaches. Furthermore, when compared to zero-shot CoT, our method exhibits significantly stronger reasoning capabilities, highlighting that understanding complex traffic scenes requires more than just the inherent knowledge of LLMs. Instead, it necessitates a systematically designed framework to interpret and reason about traffic interaction graphs. Additionally, the integration of the Frenet coordinate system proves to be a critical component, enabling more precise spatial reasoning and accurate generation.
Figure 3: Visualization of constraints. Compared to reaching the same endpoint, our method provides a more reasonable constraint, making it easier for the diffusion to follow.
Evaluation on success rate of guided simulation. To further validate the correctness of our LLM generation, we conduct a guidance success rate experiment across four types of scenarios, comparing our approach with a baseline method based on target points. The results, as shown in Table IV, demonstrate that our method achieves a significantly higher success rate than the target point-based baseline. This improvement can be attributed to the design of our Frenet-based cost functions, which provide reasonable and flexible constraints, as illustrated in Figure 3.
Figure 4: A cut-in example of refinement: Two vehicles collide with each other and the longitudinal loss term is not effective. We should increase the safety distance in the cost function.
(a) User query: vehicle 1 cuts in vehicle 2.
Figure 5: GPT solutions and visualizations for typical cases.
While our method demonstrates a high success rate, we observe that in some cases, the generated cost functions may contain errors or prove challenging for the diffusion model to follow due to different contexts in traffic scenarios. To address this, our refinement module plays a crucial role in improving the success rate, as evidenced by the results in Table IV. This module allows the LLM to adaptively adjust the cost functions, ensuring they better align with the intended guidance. Visualization examples in Figure 4 further illustrate this process, showing how the refinement module effectively corrects and optimizes the cost functions, leading to a higher success rate and more accurate guidance.
V-D Visualizations
@@
We further combine some descriptions then query the LLM, and visualize the generation results, as shown in Fig. 5. The LLM effectively decomposed the complex descriptions and generated the correct forms of cost functions. The visualizations further demonstrate that the model accurately follows the guidance of the generated functions.
Furthermore, our method demonstrates zero-shot capability in handling accident events, despite being trained exclusively on a nearly collision-free dataset, as illustrated in Fig. 6. This ability is particularly valuable for traffic simulation, as it enables the model to generalize to rare but critical scenarios, such as accidents, without requiring explicit training data.
Figure 6: Visualizations of our generated accidents.
@@ VI Conclusion
In this paper, we proposed a novel LLM-enhanced controllable traffic simulation model. Our method is capable of taking users’ descriptions as input, and generating a wide range of scenarios, including complex interactions and long-tail events. Several examples, including complex interactions and traffic accidents, have validated the effectiveness of our method. Finally, our numerous experiments and visualization demonstrate our framework achieves high-quality trajectory generation and precise instruction alignments across various benchmarks and settings.
Limitations. Although the refinement module effectively adjusts cost functions, it struggles to correct them when the LLM’s initial understanding of the scenario is fundamentally incorrect. Future work could focus on enhancing the refinement module by incorporating more robust error detection mechanisms or leveraging external knowledge sources to improve its ability to correct fundamentally flawed initial assumptions.